to solve this you can use any method for example multiplication with elimination method
using that method what is the answer to
3(x-y)=15
2x+y=7
thank you

Answer by stanbon(75887)

3(x-y)=15
2x+y=7
------
x-y = 5
2x+y = 7
----
Add to get:
3x = 12
x = 4
------
Solve for "y" using x-y = 5
---
4 -y = 5
y = -1
=============
Cheers,
Stan H.
================